Bibliography

unknown

23/03-04 16:20

[bishop1998a] Judy Bishop. Java Gently. International Computer Science Series. Addison-Wesley, 2 edition, 1998.

[bobrow1983a] Daniel G. Bobrow, Linda G. DeMichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. The LOOPS Manual. Xerox Parc, 1983.

[bosch1999a] Peter Bosch. Mixed-Media File System. PhD thesis, University of Twente, 1999.

[guttag1977b] John V. Guttag, Ellis Horowitz, and David R. Musser. Some extensions to algebraic specifications. In Language Design for Reliable Software, March 1977.

[guttag1978c] J. V. Guttag, J. J. Horning, and David R. Musser. The design of data type specifications. In Raymond T. Yeh, editor, Current Trends in Programming Methodology, Volume IV, Data Structuring, chapter 4, pages 60–79. Prentice Hall, 1978.

[itu1998a] ITU. Recommendation H.263 (02/98) – video coding for low bit rate communication. ITU-T H.263, International Telecommunication Union, 1998.

[ledoux1997a] T. Ledoux. Implementing proxy objects in a reflective ORB. In Procceings of the ECOOP’97 Workshop on CORBA: Implementation, Use and Evaluation, Jyväskylä, 1997.

[liberty2001a] Jesse Liberty. Programming C#. A Nutshell Handbook. O’Reilly & Associates, Inc., July 2001.

[liberty2001b] Jesse Liberty. Attributes and reflection. In Programming C#, A Nutshell Handbook, chapter 18. O’Reilly & Associates, Inc., July 2001.

[liebl1993a] A. Liebl. Authentication in distributed systems: A bibliography. ACM Operating Systems Review, 27(4):31–41, October 1993.

[manola1993a] F. Manola. Meta object protocol concepts for a “RISC” object model. Technical Report TR-0244-12-93-165, GTE Laboratories, 1993.

[mcilroy1968a] M. D. McIlroy. Mass produced software components. In P. Naur and B. Randell, editors, Proceedings, NATO Conference on Software Engineering, Garmisch, Germany, October 1968.

[omg1995a] Object Managment Group. The common object request broker: Architecture and specification. Technical Report 96.3.4, Object Managment Group, July 1995. (revision 2.0).

[ritchie1979a] Dennis M. Ritchie. The evolution of the unix time-sharing system. In Language Design and Programming Methodology, volume 79 of Lecture Notes in Computer Science, Sydney, Australia, September 1979. Springer-Verlag.

[tucker1997a] Allen B. Tucker, editor. The Computer Science and Engineering Handbook. CRC Press, 1997.

[weinreb1981a] D. Weinreb and D. Moon. Lisp Machine Manual. Symbolics, 1981.